OpenStudy (goformit100):

At a party, everyone shook hands with everybody else. There were 66 handshakes. How many people were at the party?

OpenStudy (goformit100):

@dmezzullo @LoveYou*69

OpenStudy (amistre64):

assume there are like 3 or 4 and see if you can develop a pattern

OpenStudy (goformit100):

i did this question several time but failed to understand the logic

OpenStudy (amistre64):

3 people, each shake 2 hands; 3*2=6, but theres duplication in that ab ba ca ac bc cb in reality acuality, ab bc ca is all

OpenStudy (amistre64):

4 people, 3 hands each ... 12 with duplication (or permutations they are called)

OpenStudy (amistre64):

do we know a way to ignore order and just count groups?

OpenStudy (shamim):

ya it will b combination

OpenStudy (amistre64):

each handshake consists of 2 people, and there are n people at the party n choose 2

OpenStudy (shamim):

ya

OpenStudy (shamim):

i know the result

OpenStudy (amistre64):

or this simplifies to:\[\frac{n(n-1)}{2}\]

OpenStudy (shamim):

=66

OpenStudy (goformit100):

Thankyou sir

OpenStudy (amistre64):

good luck ;)

OpenStudy (shamim):

anyway i think u know the law of combination is\[c _{r}^{n}=\frac{ n !}{ r!(n-r)! }\]

OpenStudy (shamim):

or\[66=\frac{ n! }{ 2!(n-2)! }\]

OpenStudy (shamim):

i think u got it